Prove that lim(Xn) =0 iff lim(|Xn|)=0 . Give an example to show that convergence of (|Xn|) need not imply the convergence of (Xn)
1
Expert's answer
2020-04-24T18:37:10-0400

"\lim\limits_{n\to\infty}x_n=0\Leftrightarrow\forall\varepsilon>0\ \exists N\ \forall n>N\ |x_n-0|<\varepsilon"

"\lim\limits_{n\to\infty}|x_n|=0\Leftrightarrow\forall\varepsilon>0\ \exists N\ \forall n>N\ ||x_n|-0|<\varepsilon"

Since "|x_n-0|=||x_n|-0|", we obtain that "\lim\limits_{n\to\infty}x_n=0\Leftrightarrow\lim\limits_{n\to\infty}|x_n|=0"

Example of a divergent sequence "x_n" such that "|x_n|" is a convergent sequence: "x_n=(-1)^n"